A taxi charges P30 for the first half kilometer and P10 for each additional 200 meters. How far in kilometers can a person ride and owe between P130 and P280?
Answer by stanbon(75887)

A taxi charges P30 for the first half kilometer and P10 for each additional 200 meters. How far in kilometers can a person ride and owe between P130 and P280?
-----------
p30 + p10(x) = P130
p10(x) = P100
x = 10
Total distance for P130 = (1/2)km + 10(200)meters = (1/2)+2 = 2.5km
-------
P30 + p(10)x = P280
p(10)x = P250
x = 25
Total distance for P280 = (1/2)km + 25(200)meters = (1/2) + 5 = 5.5km
----
Ans: He can ride between 2.5km and 5.5km
====================
Cheers,
Stan H.
====================